OpenStudy (anonymous):

Find the distance between these points. W(-6, -8), X(6, 8) 20 10 √8

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

sqrt((-6-6)^2+(-8-8)^2)=sqrt(36+64)=10

OpenStudy (anonymous):

thnx